Harry's Inc is considering a project that has the following cash flow and WACC data. What is the project's NPV? Note that if a project's projected NPV is negative, itshould be rejected-WACC 10.25% Year 0 Cash flow -1000
1 2 3 4 5
300 300 300 300 300
Answer #1


Harry's Inc is considering a project that has the following cash flow and WACC data. What is the project's NPV? Note that if a project's projected NPVis negative, it should be rejected-WACC 10.25% Year 0 Cash flow -1000
1 2 3 4 5
300 300 300 300 300


TIME

0

1

2

3

4

5

CF

-1000

PV OF CF

WACC 10.25%

PV OF CF= CF (1+RATE )^ TIME

=300 / (1+ 10.25)^1




project's NPV IS POSITIVE ACEPT THE PROJECT ---------- $130.01


answered by: BandR
Answer #2

SOLUTION ;


r = WACC = 10.25% = 0.1025 

=> 1 + r = 1.1025


NPV 

= - 1000 + 300/1.1025 + 300/1.1025^2 + 300/1.1025^3 + 300/1.1025^4 + 300/1.1025^5

= 130.01 ($) 


So, Project’s NPV is $130.01 (ANSWER)


As NPV is positive, project can be accepted. 


answered by: Tulsiram Garg